OpenStudy (anonymous):

Find the distance between the two points: Using the distance formula (-20, -4) and (-7, -6) A. 4 B. SQRT(61) C.SQRT(290) D. 5 E. SQRT(173) F. SQRT(180)

hartnn (hartnn):

Distance between points (x1,y1) and (x2,y2) is \(\huge d=\sqrt{(x_1-x_2)^2+(y_1-y_2)^2}\) does this help ??

OpenStudy (anonymous):

I already did that, and tried to solve it but i ended up getting 165 and idk how

hartnn (hartnn):

ohh... then show your work, ill find out the error .....

hartnn (hartnn):

if u used formula, tell me what u got as x1-x2 and y1-y2 = ?

OpenStudy (anonymous):

(-7 - -20)sqrt 2 +(-6 - -4)sqrt2

hartnn (hartnn):

that step is correct, what was your next step of simplification ?

hartnn (hartnn):

if its \(\large \sqrt{13^2+2^2}\) then its correct.

OpenStudy (anonymous):

ohh thats where i went wrong I had -2

OpenStudy (anonymous):

173 :D

hartnn (hartnn):

yes, sqrt 173 is correct and at first you actually get -2 but its square becomes +4 \(\sqrt{13^2+(-2)^2}=\sqrt{13^2+(2)^2}=\sqrt{173}\)